For the equation X2-13x=30, I got the answer (x-15)(X+2)by setting to zero so that: X squared -13x=30 becomes X squared -13x-30=0.
I found that -15 times 2 equals 30, and -15 plus 2 = -13. when I checked the answer, it said that X= -2 and X+15. Why isn't it 2 and -15?
Help, please?

Found 2 solutions by stanbon, Cromlix:
Answer by stanbon(75887)

For the equation X2-13x=30, I got the answer (x-15)(X+2)by setting to zero so that: X squared -13x=30 becomes X squared -13x-30=0.
I found that -15 times 2 equals 30, and -15 plus 2 = -13. when I checked the answer, it said that X= -2 and X+15. Why isn't it 2 and -15?
-----------------------
x^2 - 13x - 30 = 0
----
You factored it so you could "see" the answer:
(x-15)(x+2) = 0
-----
That product will be zero if x = +15 or is x = -2
--------------------------------
Right:
----
Procedure::
When you have (x-15)(x+2) = 0
Solve x-15 = 0, so x = 15
Solve x+2 = 0, so x = -2
---------------------------
That way you will get the correct answers.
Cheers,
Stan H.
---------------

Answer by Cromlix(4381)

Hi there,
You are correct with your
factorisation
x^2 - 13x - 30
= (x - 15)(x + 2)
To get the values for x
you make x^2 - 13x - 30 = 0
(as you did)
So,
(x - 15)(x + 2)= 0
Take each separate bracket:-
x - 15 = 0
x = 15
x + 2 = 0
x = -2
So values for x are:-
x = -2, 15.
Hope this helps :-)